THE FIFTH TERM OF AN ARITHEMETIC SEQUENCE IS 10 & THE SUM OF THE THE FIRST 10 TERMS IS 115. FIND IT'S FIRST TERM AND IT'S COMMON DIFFERENCE.
Answer by pakhi(24)

Let the 1st term be 'a' and the common difference be 'b'.
Therefore the 5th term is a+4b = 10(given)----------------(1)
The formula for the sum of 15 terms of an A.P. is
S(10)= (10/2)(2a+9b)=115
or S(10)= 5(2a+9b)=115
or S(10)= 10a+45b =115 ----------------------------------(2)

Multiplying equation1 by 10 we have 10a+40b = 100 --------(3)
Subtracting equation3 from equation2 we have 5b= 15
So we have b=3(common difference)
Putting the value of 'b' in equation1 we have a+4*3=10
So a=10-12= -2(first term of the series)